# SAQR-QC: A Logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits

### v1

@@ 7. Related Work
Quantum Hoare logic. Researchers have developed a variety of quantum Hoare logics. (RA17, ) compared three such logics, namely (CMS, ; 10.1007/978-3-642-10622-4_7, ; Ying11, ). We can divide quantum Hoare logic into expectation-based and satisfaction-based approaches. Following the approach in the seminal paper by D’Hondt and Panangaden (DP06, ), the expectation-based approaches in (Ying11, ; BHY19, ; li2019quantum, ; feng2021quantum, ) take positive operators as assertions for quantum states. This approach enables the expectation that a quantum state $\rho$ satisfies an assertion $M$ to be defined as ${\mathrm{Tr}}(M\rho)$. In contrast, the satisfaction-based logics (ZYY19, ; Unruh, ; Unruh2, ) regard subspaces of the Hilbert space as assertions. This approach enables the assertion that a quantum state $\rho$ satisfies an assertion $P$ to be defined as the support (the image space of linear operators) of $\rho$ is included in $P$. All the mentioned papers represent a predicate $M$ or $P$ as a $2^{n}\times 2^{n}$ dimensional matrix. In other words, they do not provide an efficiently computable quantum logic.
Quantum Separation logic. Quantum separation logic in (10.1109/LICS52264.2021.9470673, ) enables local reasoning for quantum computation using a quantum interpretation of Bunched Implications (o'hearn_pym_1999, ). The quantum separation logic in (10.1145/3498697, ) supports classical variables and quantum qubits’ dynamic allocation/deallocation. In these works, the separating conjunction is defined as a tensor product, i.e., quantum independence. This requirement significantly restricts the applicability of these logics.
@@
Symbolic abstraction, strongest consequence, and weakest sufficient condition The inexpressibility issues discussed in § 4.3 are a manifestation of the constraints that one faces when working with an “impoverished” logic (or logic fragment). These issues have been studied in the context of abstract interpretation as what is (now) called the symbolic-abstraction problem (DBLP:conf/vmcai/RepsSY04, ; DBLP:conf/cav/ThakurR12, ), and phrased in purely logical terms as the strongest-consequence problem (DBLP:conf/vmcai/RepsT16, , §5), as follows: {mdframed}[innerleftmargin = 3pt, innerrightmargin = 3pt, skipbelow=-0.5em] Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the strongest formula $\psi\in\mathcal{L}^{\prime}$ such that $\varphi\vDash\psi$. The strongest-consequence problem naturally arises in approximate forwards reasoning, to over-approximate a postcondition. The discussion in § 4.3 concerned backwards reasoning for which one faces the dual problem, the weakest sufficient-condition problem: {mdframed}[innerleftmargin = 3pt, innerrightmargin = 3pt, skipbelow=-0.5em] Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the weakest formula $\chi\in\mathcal{L}^{\prime}$ such that $\chi\vDash\varphi$.
As we saw in § 4.3, the strongest consequence or weakest sufficient condition may not always be expressible in $\mathcal{L}^{\prime}$, in which case one has to fall back on finding what one hopes is a suitably strong consequence or a suitably weak sufficient condition, respectively. Scherpelz et al. (DBLP:conf/pldi/ScherpelzLC07, ) presented a best-effort method for computing sufficient conditions as part of an algorithm for creating abstract transformers for use with parameterized predicate abstraction (DBLP:conf/birthday/Cousot03, ). Their method performs weakest liberal precondtion (WLP) of a post-state predicate with respect to a concrete transformer $\tau$, and then uses heuristics to identify combinations of pre-state predicates that entail the WLP value.
Reasoning about Shor’s factoring.
The QFT circuit contains many non-Clifford gates, so variants of the Gottesman-Knill theorem do not directly apply. Shor’s algorithm has been formally verified in (feng2021quantum, ; https://doi.org/10.48550/arxiv.2204.07112, ). The former uses an expectation-based approach with classical variables, while the latter employs the Coq proof assistant and the Small Quantum Intermediate Representation (sQIR) (Hicks21, ). Since Coq operates symbolically, the reasoning about quantum phase estimation in (https://doi.org/10.48550/arxiv.2204.07112, ) is symbolic.
@@
Packing of variables in abstract domains The motivation for using local observables and local projections defined with respect to a tuple $(s_{1},\cdots,s_{m})$ of sets of qubit indexes is to make reasoning scalable: each reasoning step involves only a small number of qubits. This idea is similar to the idea of “packing” variables in numeric domains, as used in Astrée (DBLP:conf/pldi/BlanchetCCFMMMR03, ): a program’s numeric variables are placed in sets (“packs”) so that each abstract transformer of a numeric abstract domain can operate on a single pack at a time. As with our qubit sets, each variable can be in multiple packs.
@@ 8. Conclusion and Future work
This paper introduces a quantitative local-reasoning framework for quantum circuits, bridging a gap in quantum program analysis. By tracking tuples of reduced-density matrices through linear functions, our method enables efficient verification of quantum circuits, including Quantum Phase Estimation, Quantum Fourier Transform, and GHZ circuits with non-Clifford gates. Unlike existing scalable frameworks limited to Clifford circuits (GottesmanKnill, ), ours extends to non-Clifford gates, significantly broadening applicability. Below, we outline future research directions for efficient reasoning about quantum programs.
One direction is to extend the logic to include multiple tuples of local observables. Specifically, we consider judgments of the form:
(32) $\{(\mathscr{A}_{1},\cdots,\mathscr{A}_{k})\,|\,\mathcal{P}\}\ \mathbf{C}\ \{(\mathscr{B}_{1},\cdots,\mathscr{B}_{k})\,|\,\mathcal{Q}\},$
where each $\mathscr{A}_{i}$ and $\mathscr{B}_{i}$ is a tuple of observables. In this setting, rather than a single inequality, a judgment yields a collection of inequalities, enabling more fine-grained and expressive reasoning about the program’s behavior. We have seen the power of this idea in the GHZ example in § 5.
It would be interesting to extend our methods to general quantum programs with classical control, such as those described in (Ying11, ; ZYY19, ). Recent papers (10.1145/3453483.3454029, ; Wu19, ; ZYY19, ) have demonstrated how to use logic to reason about noise in quantum programs. We leave the design of an efficient logic for noisy quantum programs to future work.
It is also highly compelling to explore the automation of our methods for reasoning about quantum programs. One major obstacle, shared with the automation of QAI, is the need for effective heuristics to automatically select appropriate qubit domains for analysis. An additional challenge arises from the need to resolve matrix inequalities involving positive semidefiniteness. While semidefinite constraints are common in optimization, our setting differs significantly: rather than optimizing a scalar objective function, we aim to synthesize a matrix—specifically, a postcondition observable or predicate—that satisfies a semidefinite inequality.
This requirement introduces structural difficulties. The set of positive semidefinite matrices is not a lattice under the Löwner partial order: given two such matrices $A$ and $B$, a least upper bound (i.e., a “maximum” matrix) may not exist within the set. Consequently, unlike in classical predicate synthesis or optimization, we cannot rely on lattice-theoretic fixed-point techniques to guide the construction of such matrices. Addressing this issue is essential for developing scalable and principled automated reasoning tools for quantum programs.
@@ Appendix A Appendix A: Lemmas and Proofs of Lemmas
2.1. Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$,
@@
$\rho=\sum_{i,j}\rho_{i,j}\otimes\ket{i}\bra{j},$
where $\rho_{i,j}$ are operators on subsystem $s$, and {|$\{\ket{i}\}$ is an orthonormal basis of subsystem $[n]\setminus s$. Then:
${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)$ $={\mathrm{Tr}}\left(\sum_{i,j}(A_{s}\rho_{i,j})\otimes\ket{i}\bra{j}\right)$
@@
Appendix B Appendix B: Proof of Theorem 4.2
@@
Appendix C Appendix C: Equation (20) implies Equation (15)
@@
where in going from Equation 20 to Equation 33, we used A.2; and in going from Equation 33 to Equation 34, we used Equation 19 and A.1.
Appendix D Appendix D: Reasoning about QFT using QAI
@@
We use the $\psi=|\psi\rangle\langle\psi|$ for pure state |$\ket{\psi}$,$0.x_{1}x_{2}\cdots x_{n}=\sum_{i=1}^{n}\frac{x_{i}}{2^{i}}$, and the following notation
(36) |$\ket{\psi_{x}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{2\pi i0.x}\ket{1}).$
We choose the domain ($(\{1\},\{2\},\cdots,\{n\})$, and the precondition to be
@@
where |$\ket{\psi_{j_{1}}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{2\pi i0.j_{1}}\ket{1})$ and $\psi_{j_{1}}=|\psi_{j_{1}}\rangle\langle\psi_{j_{1}}|$, by notifying $e^{2\pi i0.j_{1}}=-1$ if $j_{1}=0$, otherwise, $e^{2\pi i0.j_{1}}=1$.
@@
We continue applying the controlled-$R_{3}$, $R_{4}$ through $R_{n}$ gates and compute our postcondition, each of which adds an an extra bit to the phase of the coefficient of the first |$\ket{1}$. At the end of this procedure, we have the postcondition
@@
It follows the Swap operation between the qubit $i$ and the qubit $n+1-i$ for $1\leq i\leq n$. After applying $SWAP(1,n)$, we obtain

### v3

@@ 6. Related Work
Combining Logic and Abstract Interpretation. Combining Hoare logic with abstract interpretation is a well-established strategy in classical program analysis. Many software model-checking tools integrate logic-driven state-space exploration with dataflow information obtained via abstract interpretation, often computed as a preliminary pass. A prominent example is the Seahorn solver for Constrained Horn Clauses, whose Crab component provides abstract-interpretation capabilities (Crab, ; GN21, ). This integration illustrates how abstract interpretation can complement logical reasoning, resulting in verification that is both scalable and precise. Our work extends this principle to the quantum domain, where the combination supports reasoning about program semantics that would be difficult to achieve with either technique in isolation.
Quantum Hoare Logic. Researchers have developed a variety of quantum Hoare logics. A comparative study in (RA17, ) examined three representative logics (CMS, ; 10.1007/978-3-642-10622-4_7, ; Ying11, ). Quantum Hoare logics can be broadly categorized into expectation-based and satisfaction-based approaches. Following the seminal work of D’Hondt and Panangaden (DP06, ), expectation-based approaches (Ying11, ; BHY19, ; li2019quantum, ; feng2021quantum, ) use positive operators as assertions for quantum states, defining the expectation that a state $\rho$ satisfies an assertion $M$ as ${\mathrm{Tr}}(M\rho)$. In contrast, satisfaction-based logics—exemplified by Zhou et al. (ZYY19, ) and indepedently by Unruh (Unruh, ; Unruh2, )—treat subspaces of the Hilbert space as assertions, providing a direct semantic correspondence between program correctness and the geometry of quantum states. Here, a state $\rho$ satisfies an assertion $P$ if the support of $\rho$ is included in $P$. In all of these works, predicates $M$ or $P$ are represented as $2^{n}\times 2^{n}$ matrices, which precludes efficient computation and limits scalability for larger quantum systems.
Quantum Separation logic. Quantum reasoning frameworks have made significant progress along multiple fronts. Quantum separation logic, introduced in (zhou2021quantum, ), extends the principles of Bunched Implications (o'hearn_pym_1999, ) to the quantum setting, enabling local reasoning about quantum programs. The framework in (10.1145/3498697, ) further incorporates hybrid state spaces by supporting both classical variables and the dynamic allocation and deallocation of quantum qubits. However, in both approaches, the separating conjunction is interpreted strictly as a tensor product—corresponding to quantum independence—which substantially limits the expressiveness and applicability of the logic. In addition, because predicates are restricted to projection operators, these systems support only qualitative rather than quantitative reasoning about quantum states and their interactions.
@@
Symbolic abstraction, strongest consequence, and weakest sufficient condition. The inexpressibility issues discussed in §˜3.2 are a manifestation of the constraints that one faces when working with an “impoverished” logic (or logic fragment). These issues have been studied in the context of abstract interpretation as what is (now) called the symbolic-abstraction problem (DBLP:conf/vmcai/RepsSY04, ; DBLP:conf/cav/ThakurR12, ), and phrased in purely logical terms as the strongest-consequence problem (DBLP:conf/vmcai/RepsT16, , §5), as follows: Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the strongest formula $\psi\in\mathcal{L}^{\prime}$ such that $\varphi\vDash\psi$. The strongest-consequence problem naturally arises in approximate forwards reasoning, to over-approximate a postcondition. The discussion in §˜3.2 concerned backwards reasoning for which one faces the dual problem, the weakest sufficient-condition problem: Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the weakest formula $\chi\in\mathcal{L}^{\prime}$ such that $\chi\vDash\varphi$.
As we saw in §˜3.2, the strongest consequence or weakest sufficient condition may not always be expressible in $\mathcal{L}^{\prime}$, in which case one has to fall back on finding what one hopes is a suitably strong consequence or a suitably weak sufficient condition, respectively. Scherpelz et al. (DBLP:conf/pldi/ScherpelzLC07, ) presented a best-effort method for computing sufficient conditions as part of an algorithm for creating abstract transformers for use with parameterized predicate abstraction (DBLP:conf/birthday/Cousot03, ). Their method performs weakest liberal precondtion (WLP) of a post-state predicate with respect to a concrete transformer $\tau$, and then uses heuristics to identify combinations of pre-state predicates that entail the WLP value.
Reasoning about Shor’s factoring. The QFT circuit contains many non-Clifford gates, so variants of the Gottesman-Knill theorem do not directly apply. Shor’s algorithm has been formally verified in (feng2021quantum, ; https://doi.org/10.48550/arxiv.2204.07112, ). The former uses an expectation-based approach with classical variables, while the latter employs the Coq proof assistant and the Small Quantum Intermediate Representation (sQIR) (Hicks21, ). Since Coq operates symbolically, the reasoning about quantum phase estimation in (https://doi.org/10.48550/arxiv.2204.07112, ) is symbolic.
@@
Packing of variables in abstract domains. The motivation for using local observables and local projections defined with respect to a tuple $(s_{1},\cdots,s_{m})$ of sets of qubit indexes is to make reasoning scalable: each reasoning step involves only a small number of qubits. This idea is similar to the idea of “packing” variables in numeric domains, as used in Astrée (DBLP:conf/pldi/BlanchetCCFMMMR03, ): a program’s numeric variables are placed in sets (“packs”) so that each abstract transformer of a numeric abstract domain can operate on a single pack at a time. As with our qubit sets, each variable can be in multiple packs.
@@ 7. Conclusion and Future work
This paper introduces a framework for local reasoning about quantum circuits that is both scalable and addresses quantitative properties. The framework combines Quantum Hoare Logic, which tracks quantitative properties of quantum states, with quantum abstract interpretation, which provides scalable reasoning about qualitative properties. Unlike a prior scalable framework that is restricted to Clifford circuits (GottesmanKnill, ), our approach supports non-Clifford gates, greatly extending its applicability. By tracking tuples of reduced density matrices via tuples of local observables, it fills a long-standing gap in methods for reasoning about quantum programs, enabling scalable and quantitative verification of quantum circuits on classical computers.
We apply our logic to GHZ circuits with non-Clifford unitaries, successfully characterizing the output state up to a relative phase. We further demonstrate its use on the Quantum Fourier Transform, and derive a lower bound on the success probability of the Quantum Phase Estimation algorithm.
Extending our methods to general quantum programs with classical control (Ying11, ; zhou2021quantum, ) and designing a logic for noisy programs (10.1145/3453483.3454029, ; Wu19, ; ZYY19, ) are promising directions for future work. Another direction, automating our reasoning framework, will require solving two fundamental challenges: (i) how to select qubit domains, and (ii) how to synthesize predicates as positive semidefinite observables under semidefinite constraints.
@@ Appendix A Appendix A: Lemmas and Proofs of Lemmas
˜2.1. Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$,
@@
$|,\rho=\sum_{i,j}\rho_{i,j}\otimes\ket{i}\bra{j}$,
[…]
@@
Appendix B Appendix C: Proof of Theorem 3.1
Proof.
Assume that $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ as defined in ˜3.8. Then for any $\rho$, we have
Let us choose $\rho\models\mathcal{P}$. According to Lemma 2.5, we have
On the other hand,
${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$
$=$ $\sum_{i}1=m.$
where the inequality follows from ˜2.4 and $Q_{s_{i}}\leq I_{s_{i}}$.
Therefore, ${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$ and ${\mathrm{Tr}}(Q_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))=1$ for all ii. According to Lemma 2.5, we know ${\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho)\vDash Q_{s_{i}}$, which implies
$\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\},$
where LABEL:qai $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$ iff for all $\rho$, $\rho\vDash\mathcal{P}$ implies $\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q}$. ∎
Appendix C Appendix C: Proof of Theorem 3.2
@@
Appendix D Appendix D: Equation (16) implies Equation (11)
@@
where in going from Equation˜16 to Equation˜27, we used ˜A.2; and in going from Equation˜27 to Equation˜28, we used Equation˜15 and ˜A.1.
Appendix E Appendix E: Reasoning about QFT using QAI
@@
[…]
We choose the domain $(\{1\},\{2\},\cdots,\{n\})$, and the precondition to be
@@
[…]
@@
[…]
@@
It follows the Swap operation between the qubit $i$ and the qubit n+1in+1-i for $1\leq i\leq n$. After applying $SWAP(1,n)$, we obtain
@@
Appendix F Appendix F: Proofs of Eq. (24) and Eq. (26) in Section 5
F.1. Proof of Equation˜24
Equation˜24 is equivalent to
To see this, we observe
Let $V=H(1)^{-1}CR_{2}(2,1)^{-1}CR_{3}(3,1)^{-1}\cdots H(n-k+2)^{-1}$ means
$U_{k}=H(n-k+1)^{-1}\cdots H(n-1)^{-1}CR_{2}(n,n-1)^{-1}H(n)^{-1},$
Then
This argument proves Equation˜24.
We first use quantum abstract interpretation (YP21, ) to prove
After applying the first $H$ gates, the post-condition becomes
Similarly for $CU^{2^{n-k+1}}\cdots CU^{2^{0}}$. This argument proves Equation˜31.
F.2. Proof of Equation˜26
To prove Equation˜26, we observe
and
where
$V(C_{1})$ $=CU^{2^{n-k}}\cdots CU^{2^{n-1}}$
[…]
$=|\langle\tau|\omega\rangle|^{2}$
[…]
$=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]$
$=r.$